Kathy is 20 years older than Carol. In 6 years, the sum of their ages will be 50. How old is each now?
Answer by josgarithmetic(39799)

k and c for kathy's age and carol's age NOW.

Now: k=c+20, which becomes k-c=20.
Six Years from Now: (k+6)+(c+6)=50, from which
k+c=50-12,
k+c=38.

You have this system to solve:
k-c=20, and k+c=38.
Various few simple ways to continue. Try k=c+20, one of the original equations, and substitute into k+c=38..., solve for c...